food goblin supermarkets use both cashiers and baggers to serve customers at check out. during the first 7 hours of each workday (monday–friday), 5 cashiers and 4 baggers serve approximately 24 customers per hour. a cashier and a bagger who require approximately 4 minutes at checkout and 2 minutes at bagging serve each customer.

a. How many employees should food goblin supermarket schedule if it requires a 8 percent capacity cushion?

Food goblin supermarket should schedule( 2 ) cashier(s)( Enter your response rounded up to the nearest whole number)

Food globin supermarket should schedule ( 1) bagger(s) ( Enter your response rounded up to the nearest whole number)

B. Assume that both baggers and cashiers are cross-trained to perform both activities so that they can serve customers independently. Customers are now both checked and their groceries bagged by one individual. Further, assume that it takes 7 minutes for one individual to both cash out and bag each customer’s groceries. How many employees should food goblin supermarket schedule if it requires as 8 percent capacity cushion?

Food Goblin supermarket should schedule( ) cross-trained employees. (Enter your response rounded up to the nearest whole number.)
Answer #1

Total customers that can be served alone by 5cashiers per hour = 60*5 /4 = 75 customers

Utilization of cashiers = Output / Maximum capacity = 24 customers per hour /75= 32%

Total customers that can be served alone by 4 baggers per hour = 60*4 /2 = 120 customers

Process capacity = 75 customers / hour (Bottleneck)

Utilization of baggers = 24/120 = 20%

Cashiers utilization should be Current Capacity + Capacity cushion =32+8 = 40%

For 40% utilization Customers served = 40% * 75 = 30 customers per hour

Cashier required for 30 customers = 30/15 = 2 Cashier (Note - One cashier can server 60/4 = 15 customers per hour)

Similarly utilization of baggers required = 20+8% = 28%

For 28% utilization Customers served = 28% * 75 = 21

One bagger can serve 30 customers in an hour so Baggers required = 1

part b) Process capacity per employee per hour = 60 minutes / 7 = 60 / 7 per hour

To seve 24 customers in an hour employees required = 24*7 / (60) = 2.8 employees (100% utilization)

However, we need to account 92% utilization, considering 8% capacity cushion, Cross-trained Employees required = 2.8 * 100 /92 = 3.04 or 3